Operations Report – December 11th

Crew 200 Operations Report 11-Dec-2018

SOL: 9

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart
Notes on non-nominal systems: None

Generator (hours run): From 4:50 pm last night to 8:10 am. On tonight at 5 pm.

Solar— SOC 78%
Diesel Reading – 24 %
Propane Reading – 31 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 350 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – high level
Water (loft) – Static to Loft Pump used – yes
Water Meter: 2

Toilet tank emptied: no

Deimos rover used: no, still not functional
Hours:141.5
Beginning charge:
Ending charge: 71%
Currently charging: no

Sojourner rover used: No
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: no
Hours: 59.6
Beginning charge:
Ending charge: 100%
Currently charging: yes

Opportunity rover used: yes
Hours: 41.4
Beginning charge: 100%
Ending charge: 67%
Currently charging: yes

Curiosity rover used: yes
Hours: 57.9
Beginning charge: 100%
Ending charge: 57%
Currently charging: yes

Notes on rovers: We were a bit surprised by how fast Curiosity rover’s battery was going down (down to 57% after 30mn)
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: Eveything looks fine

Summary of internet: Nothing to report.

Summary of suits and radios: we performed an EVA this morning to study the radio coverage, everything was alright. Ventilation of spacesuit 3 still not functional (I tried with another charger)

Summary of Hab operations: everything is fine, living, cooking and sleeping are still pleasant.

Summary of GreenHab operations: Regular activities

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Nominal, stayed closed during the day, is now open.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

EVA Report – December 11th

Crew 200 EVA Report 11-12-2018

EVA #4

Author: Antoine Bocquier

Purpose of EVA: 1) Map out locations that allow radio communication between EVA crew and HabCom, in northern and southern regions 2) Test how EVAs can safely and effectively be conducted with three-member crew, including rover use.

Start time: 9:10

End time: 10:55

Narrative: After the engineering tour, we drove with Curiosity to the north, then to the south to explore and map the regions of communication strength/readibility. We applied a communication protocol to control the radio strength/readability (QSA/QRK codes) every minute and communicate GPS coordinates every 5mn.
We first drove to Galileo road, planning to stop for a short walk at Connector Stream Bed. However we did not find both of them, being focused on our communication test protocol and with the recent snow melting. We went further than expected to the north, up to cowboy’s corner where we stopped as we realized from the transmitted coordinates that we were too far.
On our way back to the Hab to change the rover and go to the south, we confirmed the location of Connector Stream Bed. Arrived at the Hab at 9:50, we took Opportunity to go to Robert’s Rock Garden, applying the same communication protocol. From there we walked along Kissing Camel Ridge Edge to complete yesterday’s mapping of this area. We then drove back and came back on time at the Hab. Analysis of the acquired data is in process and the first results are promising!

Destination: Galileo Road (up to Cowboy Corner) + Robert’s Rock Garden

Coordinates (use UTM NAD27 CONUS): 518000 E – 4249300 N to 519000 E – 4252500 N

Participants: Makiah, Antoine

Road(s) and routes per MDRS Map : Cow Dung Road by driving, walking along Kissing Camel East Ridge

Mode of travel: driving, walking

Sol Summary – December 11th

Sol: 9

Summary Title: First 3 crew EVA using rovers.

Author’s name: Lindsay Rutter

Mission Status: All crew members are happy after 3 crew EVA using rovers.

Sol Activity Summary: We conducted a 3 crew EVA today using rovers and
continued mapping communication strength and readibility. We realized
yesterday that we did not document any photos during our first 3 crew
EVA because we were concentrating on our communication protocol.
Today, we reserved extra time at the beginning and end of our EVA to
better document it with photos. After our EVA, we added our newest
observations about 3 crew EVAs to a list we are keeping this week and
will share at the end of our mission. We continued plotting the radio
test values we collected during our EVA onto maps and collaborating
with projects in the habitat. We also caught a habitat mouse today;
tomorrow, we will remove it a significant distance from the habitat as
per protocol. We will be careful to monitor and report any other
problematic pests around the habitat.

Look Ahead Plan: We plan to continue mapping the radio connectivity
between EVA crew and HabCom in northern and southern areas and testing
how to safely and effectively conduct 3 crew EVAs. We will also
continue plotting our tested radio connectivity values onto maps
together and collaborating between projects in the habitat and
GreenHab.

Anomalies in work: Mouse was found in a trap inside the habitat. We
will follow protocol and remove it during our EVA tomorrow.

Weather: Cold, muddy from snow melt.

Crew Physical Status: No crew members reported physical problems today.

EVA: Crew completed an EVA with 2 members outside and 1 member in the
habitat. EVA members performed an engineering check, drove north with
Curiosity to Cowboys’ corner (this was further than intended due to
recent snow melting obstructing landmarks), returned to the habitat,
switched rovers, drove south with Opportunity to Robert’s Rock Garden,
walked along Kissing Camel Ridge, returned to the habitat, took
several photos to document our 3 crew EVA. We applied our
communication and GPS coordinate relay system throughout the EVA.

Reports to be filed: Operations Report, EVA Report, EVA Request,
GreenHab Report, Pictures Report, Journalist Report, Sol Report

Support Requested: N/A

Journalist Report – December 11th

Crew 200 Journalist Report 11-DEC-2018

*In spoken word format*

Where would I be if I chose education over learning?
I would have arose this morning, on Earth.
I would’ve taken the exam that kept me on Earth, yesterday.
Because all things considered, at the end of the day,
I am here for school.
A professor will rule a grade that affects my career.
My research on Mars doesn’t get graded at the end of year.
The rules say I’m a student before an engineer.

Well, this isolation does something at MDRS, especially for the
engineering mind.
Where the problem is undefined.
The method is up to you,
but the solution is experience and time.

At the end of the day,
when you take your course books offline,
your education is now experiential
Your learning has boundless potential

Heat Transfer lessons leap off binded textbooks
As Antoine finds nooks in the Hab to refine his model.
Computer Science lab is now carried
by Lindsay’s work to meet the need
of a report reading software.
Everyday, we code through another layer.

I’m on EVA, my phone on display to the GPS received from 16 of 23 satellites
I learned about in a lecture in class. I might
lose signal with the Hab Com, but we have protocol,
procedures, to anticipate all
scenarios so we can come home safe.
We did. Not a minute too soon or too late.

Do I have my priorities straight?
Would it have been better to wait until I graduate?
Or until I have Dr. in front of my name,
Where a professor would have respect for me same?
When the difference between Mars and Earth
is only a final examination,
I will choose learning over education.

Makiah Eustice

Greenhab Report – December 11th

Crew 200 Greenhab Report 11-DEC-2018

Greenhab Officer:
Makiah Eustice

Environmental Control:
Heating
40% Shade Cloth on

Average Temperatures:
Low: 24 C
High: 54 C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
8 gallons

Water in Blue Tank (lbs):
68%

Times of Watering for Crops:
1430

Changes to crops:
NONE

Narrative:
You can tell the tomatoes are trying to grow again. My sprouts will be
ready to harvest before we leave. I put up new sticky sheets for the
gnats.

Harvest:
NONE

Support/Supplies Needed:
NONE